OpenStudy (anonymous):

Find the point that splits segment EF in half if point E is located at (5, -2) and point F is located at (-1, 3). Select one: a. (1.5, 1) b. (2, 0.5) c. (3.5, 4.5) d. (3, -2.5)

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

We will need the midpoint formula. See attachment.

Directrix (directrix):

So, the x-coordinate of the midpoint is (5- 1)/2 = ? The y-coordinate of the midpoint is (-2 +3)/2 = ? Hint: Think about it as the average of the two x-coordinates and then the average of the two y-coordinates. @prowrestler

OpenStudy (anonymous):

im back @Directrix

Directrix (directrix):

Then, do this: What is this: (5- 1)/2 = 4/2 = ? That is for x.